IP You are driving through town at 12.5 m/s when suddenly a ball rolls out in front of you. You apply the brakes and begin decelerating at 3.7 m/s2. How far do you travel before stopping?
Answer #1

here,

the initial speed of car , u = 12.5 m/s

accelration of car , a = - 3.7 m /s^2

let the distance travelled before stopping be s

final speed is zero

using third equation of motion

v^2 - u^2 = 2 * a * s

0 - 12.5^2 = - 2 * 3.7 * s

s = 21.1 m

the distance travelled before stopping is 21.1 m